OpenStudy (anonymous):

The pipe in question is part of a straight pipeline water flow system. The flow within this system is due to ΔP (Pressure drop) across the length of a pipe (L=300m), carrying water, with diameter (D=0.2m), flowing at a rate of 1.7 m/s. The pipe is made out of cast iron (e=0.25mm) The following data needs to be applied: • The losses due to entrance of the pipe (k=0.5) • The losses due to exit of the pipe (k=0.3) • NRE = ρVD/μ Determine the head losses in pipeline flow; 1. due to friction (both minor and major) 2. due to entrance and exit 3. pressure head losses within the system

OpenStudy (surry99):

I can help you with this...what have you tried so far?

OpenStudy (anonymous):

I have tried calculating the friction factor by dividing 0.25mm by 200mm as the diameter is 0.2m, I am unsure where to use this friction factor figure?

OpenStudy (surry99):

The friction factor is obtained from the Moody chart. Are you familiar with that?

OpenStudy (surry99):

I have to run but will be back....

OpenStudy (surry99):

First step is to calculate the Reynolds number for the flow...will be back in a couple of hours.

OpenStudy (anonymous):

\[RE=((0.2*1.7*1000)\div0.001)=340,000 \]

OpenStudy (anonymous):

I have tried this so far but my lecturer only ever gave us one example of this equation so I cannot be sure if 340,000 is a reasonable answer?

OpenStudy (surry99):

That look ok...now go to the Moody chart and look up the friction factor using: your Reynolds number and e/d...gotta run now...will be back

OpenStudy (anonymous):

I am unsure how to use the reynolds number at this point but I think e/d = 0.00125 so it seems the figure would be 2.3 on the moody chart, does this sound correct?

OpenStudy (surry99):

So Re = 3.4 E5 e/d = .25mm/200 mm = .00125 That is correct....now the friction factor where these intersect is about .02

OpenStudy (surry99):

Let me now if you see where I got .02 for the friction factor off the Moody chart

OpenStudy (anonymous):

Yes I can see, it is on the left on my graph

OpenStudy (surry99):

Great...now you need to apply your equations for head loss for: 1) straight pipe length 2) entrance loss 3) exit loss Check your notes to see if you have those equations

OpenStudy (anonymous):

\[-f (L \div D) * ( v^2 \div 2 ) * \rho\] Can I use this equation to solve the straight pipe length?

OpenStudy (surry99):

h = f*L/D*V^2/2....where di d you get the rho from?

OpenStudy (anonymous):

Its just because I have a variant of the equation I think, it ends in (v^2/2g)*pg but the g's cancel out. Must be for something else. Anyway.... \[0.02\times(300\div0.2)\times(0.7^{2}\div2) = 7.35\] Does this seem like a reasonable figure?

OpenStudy (surry99):

V = 1.7 not 0.7 ...be careful

OpenStudy (surry99):

sorry....my bad v = 1/7 m/s = .143 m/s

OpenStudy (surry99):

Better check the problem ...is V = 1.7 m/s or 1/7 m/s??

OpenStudy (anonymous):

Sorry about that it was 1.7^2 0.02*(300/0.2)*(1.7^2/2) = 43.35

OpenStudy (surry99):

ok...just make sure the problem said 1.7 m/s

OpenStudy (surry99):

Now you have to calculate the losses at entrance and exit...were you provided with formula for that?

OpenStudy (anonymous):

Okay I just checked and it is definitely 1.7 m/s. No I wasn't given a formula for this but I have been looking at using... \[Kentrance(V ^{2}\div2g)\] And the other K factor the exit loss, Does this seem acceptable?

OpenStudy (surry99):

h loss = K*V^2/2

OpenStudy (anonymous):

That definitely looks familiar... \[k \times (v ^{2}\div(2\times9.81)) = \] Entrance 0.5 * (1.7^2/(2*9.81)) = 28.35 Exit 0.3 * (1.7^2/(2*9.81)) = 17.01 Am I doing this correctly?

OpenStudy (surry99):

I don't have g in my equation....

OpenStudy (surry99):

Check your text to see what equations you should be using...some have g

OpenStudy (anonymous):

I would rather trust your equation, all of mine include g but my lecturer has messed up on many equations

OpenStudy (anonymous):

Ok my answers are Entrance \[k=0.5\times(0.7^{2}\div2)=0.1225\] Exit \[k=0.3\times(0.7^{2}\div2)=0.0735\] Is that a better result?

OpenStudy (anonymous):

No.... I mean 1.7^2 k0.5 = 0.7225 k0.3 = 0.4335

OpenStudy (anonymous):

I guess these are the pressure head losses in the entrance and exit? Am I to designate Pa as their unit or KPa if I am to do this in S.I units?

OpenStudy (surry99):

Just got back ...give me a few minutes and we can finish this up

OpenStudy (surry99):

if your instructor is writing the Bernoulli equation in the following manner, then the head loss equations he/she provided to you are correct: P + 1/2*rho*v^2 + rho*g*h

OpenStudy (anonymous):

I was wondering whether I need to use the 'h' because it is just a straight pipe, is that correct? \[(\frac{ 1 }{ 2 }\times \rho \times) v ^{2} + (\rho \times g)\] Can I use the Bernoulli equation as shown above?

OpenStudy (surry99):

what you have above is part of the Bernoulli equation...compare it to mine to see what you are missing

OpenStudy (surry99):

Use the equations provided by your instructor and put in all the units to ensure your calculations for the head loss is in the correct units.

OpenStudy (anonymous):

The Bernoulli equation that I was provided with is shown below... \[\Delta ( \frac{ \rho }{ \rho g } + \frac{ v ^{2} }{ 2g } + z)\] But I have been told that if the pipe is straight horizontal then we do not need to use the 'z' as it is for gradient.

OpenStudy (anonymous):

I think maybe I need to use... \[( \frac{ p _{1} }{ \rho } ) + (\frac{ v ^{2} }{ 2 } )\] For entrance and... \[( \frac{ p _{2} }{ \rho } ) + (\frac{ v ^{2} }{ 2 } )\] For exit? If I use the figures generated from the K factor equations as P1 and P2 would this give me a feasible answer?